OpenStudy (anonymous):

find the exact value of sin((11π)/(6))? pretty easy

OpenStudy (paxpolaris):

ALL right ... what's your answer:

OpenStudy (anonymous):

some huge decimal... is that right that why asking .1003545349

OpenStudy (anonymous):

thats why im*

OpenStudy (anonymous):

\[ \sin \left( \frac {11\pi}6 \right)=\\ \sin \left( \frac {12\pi}6 -\frac \pi 6 \right)= \sin( 2 \pi -\frac \pi 6)=-\frac 12 \]

OpenStudy (paxpolaris):

-1/2